# Md Kowsher School of Engineering and Science

### v1

@@ 4.1 Result Analysis
Dataset Electricity 96 0.140 0.141 0.140 0.193 0.201 0.274 0.386 0.258 1.588 0.140
192 0.153 0.154 0.153 0.201 0.222 0.296 0.386 0.266 1.595 0.151
336 0.169 0.171 0.169 0.214 0.231 0.300 0.378 0.280 1.617 0.168
720 0.203 0.210 0.203 0.246 0.254 0.373 0.376 0.283 1.647 0.201
Exchange 96 0.082 0.089 0.081 0.148 0.197 0.847 0.376 0.968 0.081 0.101
192 0.167 0.180 0.157 0.271 0.300 1.204 1.748 1.040 0.167 0.192
336 0.328 0.331 0.305 0.460 0.509 1.672 1.874 1.659 0.305 0.305
720 0.964 1.033 0.643 1.195 1.447 2.478 1.943 1.941 0.823 0.586
Traffic 96 0.410 0.410 0.410 0.587 0.613 0.719 2.085 0.684 2.723 0.391
192 0.423 0.423 0.423 0.604 0.616 0.696 0.867 0.685 2.756 0.399
336 0.436 0.435 0.436 0.621 0.622 0.777 0.869 0.734 2.791 0.425
720 0.466 0.464 0.466 0.626 0.660 0.864 0.881 0.717 2.811 0.442
Weather 96 0.176 0.182 0.176 0.217 0.266 0.300 0.896 0.458 0.259 0.173
192 0.218 0.225 0.220 0.276 0.307 0.598 0.622 0.658 0.309 0.217
336 0.262 0.271 0.265 0.339 0.359 0.578 0.739 0.797 0.377 0.259
720 0.326 0.338 0.323 0.403 0.419 1.059 1.004 0.869 0.465 0.314
ILI 24 1.947 1.683 2.215 3.228 3.483 5.764 1.420 4.480 6.587 0.109
36 2.182 1.703 1.963 2.670 3.103 4.755 7.394 4.799 7.130 0.109
48 2.256 1.719 2.130 2.622 2.669 4.763 7.551 4.800 6.575 0.111
60 2.390 1.819 2.368 2.857 2.770 5.264 7.662 5.278 5.893 0.113
ETTh1 96 0.375 0.374 0.375 0.376 0.449 0.865 0.664 0.878 1.295 0.375
192 0.418 0.408 0.405 0.420 0.500 1.008 0.790 1.037 1.325 0.411
336 0.479 0.429 0.439 0.459 0.521 1.107 0.891 1.238 1.323 0.432
720 0.624 0.440 0.472 0.506 0.514 1.181 0.963 1.135 1.339 0.436
ETTh2 96 0.288 0.277 0.289 0.346 0.358 3.755 0.645 2.116 0.432 0.228
192 0.377 0.344 0.383 0.429 0.456 5.602 0.788 4.315 0.534 0.269
336 0.452 0.357 0.448 0.496 0.482 4.721 0.907 1.124 0.591 0.289
720 0.698 0.394 0.605 0.463 0.515 3.647 0.963 3.188 0.588 0.306
ETTm1 96 0.308 0.306 0.299 0.379 0.505 0.672 0.543 0.600 1.214 0.312
192 0.340 0.349 0.335 0.426 0.553 0.795 0.557 0.837 1.261 0.348
336 0.376 0.375 0.369 0.445 0.621 1.212 0.754 1.124 1.283 0.372
720 0.440 0.433 0.425 0.543 0.671 1.166 0.908 1.153 1.319 0.424
ETTm2 96 0.168 0.167 0.167 0.203 0.255 0.365 0.435 0.768 0.266 0.151
192 0.232 0.221 0.224 0.269 0.281 0.533 0.730 0.989 0.340 0.187
336 0.320 0.274 0.281 0.325 0.339 1.363 1.201 1.334 0.412 0.210
720 0.413 0.368 0.397 0.421 0.433 3.379 3.625 3.048 0.521 0.285
Table 3: Comparing multivariate long-term forecasting errors using MSE: lower values indicate better performance. Specifically, for the ILI dataset, forecasting horizons are $T\in\{24,36,48,60\}$, while for the remaining datasets, they are $T\in\{96,192,336,720\}$. Optimal results are emphasized in bold, while the top results from Transformers are underlined for distinction.”
Horizons 100 500 1500 2500
Dataset Model MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 NLinear RT 0.362 0.371 0.399 0.403 0.439 0.438 0.479 0.442 0.692 0.602 0.798 0.714 1.683 1.294 2.832 2.104
ETTh2 NLinear RT 0.283 0.223 0.350 0.310 0.684 0.303 0.591 0.574 0.947 0.561 1.822 0.935 1.361 0.951 2.891 1.821
ETTm1 NLinear RT 0.302 0.311 0.350 0.352 0.441 0.357 0.433 0.418 0.849 0.739 1.476 1.102 1.153 0.896 2.768 1.562
ETTm2 NLinear RT 0.164 0.127 0.240 0.217 0.411 0.267 0.432 0.329 1.084 0.740 2.148 0.977 2.099 1.180 3.260 1.936
Table 4: Comparison of NLinear and RT for explaining the Lyapunov Exponent’s predictive performance across different datasets. The table presents MSE and MAE values for both models at various prediction horizons: 100, 500, 1500, and 2500. The results demonstrate that the RT model consistently outperforms the NLinear model, achieving significantly lower MSE and MAE values, thus overcoming the Lyapunov Exponent’s challenges of far-sighted forecasting.
Figure 3: Visualization comparing the prediction of the Lyapunov Exponent using NLinear and RT models. The X-axis corresponds to the time sequence, while the Y-axis corresponds to the target feature.
4.2 Ablation Study
Lyapunov Exponent: The Lyapunov Exponent (LE) measures how chaotic a system is by computing how quickly points on a path move away from each other over time. Table 4 shows our RT and a baseline NLinear LE over different time steps of 100, 500, 1500, and 2500, in future on different datasets (ETTh1, ETTh2, ETTm1, ETTm2) and outperforms the NLinear model in predicting the LE. Figure 3 illustrates the comparison of forecasting 2500 future time steps between the NLinear and RT models, where RT matches more the ground truth than NLinear.
Entropy Datasets $\vec{u}(t)$ $\vec{o}(t)$ $\vec{z}(t)$ ETTh1 1.011 1.966 2.331 ETTh2 0.993 1.892 2.015 ETTm1 0.816 2.056 2.831 ETTm2 0.599 1.248 2.131 [Uncaptioned image] Table 5: Top: Entropy in different datasets. This table presents the entropy values of the original inputs ($\vec{u}(t)$), the corresponding reservoir readouts ($\vec{o}(t)$), and the resultant transformed representations ($\vec{z}(t)$) across various datasets. Bottom: The Lime Explanation effect of current input (A) and reservoir (B) on output prediction is shown for ETTh1
[Uncaptioned image] Figure 4: In the presented graph, the X-axis represents the size of the model parameters in thousands, while the Y-axis illustrates the mean absolute error of the prediction outputs. The yellow boxes in the graph indicate the history lengths used in the computation of the RT-1R, RT-5R, and RT-10R models, which were each constructed with one, five, and ten reservoirs, respectively.
Explain Signature Features: We illustrate and measure how well-distributed the features input of Transformer with or without reservoir are. More distribution means higher entropy and more informative, and vice versa. We use the Local Interpretable Model-agnostic Explanations (LIME) algorithm (Ribeiro et al., 2016). Our goal is to explain the decision-making process within the readout layer by employing LIME to generate a local linear approximation of the model. The initial step involves generating perturbed input time series by adding noise to the original $\vec{u(t)}$. A simpler interpretable model is then trained using these perturbed inputs along with corresponding output time series $\vec{y}(t)$. The complexity of the interpretable model depends on the reservoir model’s intricacy, encompassing both linear and non-linear options. Ultimately, we utilize the coefficients of the interpretable model, along with LIME-calculated feature importances, to explain the readout layer’s decision. This Table (Top) 5 shows that the readout from the reservoir, $\vec{o}(t)$, has higher entropy than Transformer original input features $\vec{u}(t)$. By concatenating both using Equation 11, we achieve the highest entropy. Additionally, Table (Bottom) 5 illustrates the Lime explanation for both RT and the baseline Transformer (more ablation in the Appendix A.4).
@@ 4.2 Ablation Study
History length versus parameter size: Conventional approaches for modeling temporal data require a large number of training parameters, proportional to the window size (s) and the number of features (C). This becomes infeasible when considering long historical sequences. In contrast, reservoir computing approaches, such as the proposed method, only train a vector representation (i.e. readout) for the next time step without explicit conditioning on the entire history. Empirical results, illustrated in Figure 4, demonstrate that RT provides a better trade-off between the number of training parameters and loss.
Impact of Reservoir Initial Conditions
Figure 5: Comparison of MSE loss across five reservoir initialization techniques: random initialization, zero initialization, constant initialization, normal distribution initialization, and uniform distribution initialization. The x-axis represents the initialization methods, while the y-axis denotes the corresponding MSE loss values.
In this study, we delved into the intricate dynamics of reservoir initialization and its consequential impact on the performance of our RT models. The reservoir initialization process lays the foundation for the system’s dynamic behavior, influencing the trajectory of internal states and subsequent information processing capabilities. We investigated the impact of different initialization conditions on the performance of our RT model. We employed five distinct initialization approaches, both in ensemble configurations with group reservoirs and without group reservoirs, as depicted in Figure 5.
Firstly, we utilized random initialization, which initializes the reservoir randomly within the range of 0 to 1. Subsequently, we employed zero initialization, which sets all reservoir values to zero. We also explored constant initialization, where a constant value is randomly selected for initialization. Furthermore, we investigated initialization using normal and uniform distributions.
Our experimentation revealed that among these initialization methods, zero initialization consistently yielded superior results in terms of MSE loss, both in ensemble setups with and without group reservoirs. Notably, the ensemble technique with group reservoirs demonstrated the lowest loss across both the ETTh2 and ETTm2 datasets.
@@ 5 Related Work
Chaotic time-series prediction using machine learning approaches has been investigated in the last decade using CNN, WNN, FNN, LSTM, etc. (Ramadevi & Bingi, 2022). For example,(Sangiorgio & Dercole, 2020) proposed the use of a Recurrent Neural Network (RNN) with Long Short-Term Memory (LSTM) cells for chaotic time-series prediction and demonstrated improved performance on several chaotic systems.(Karasu & Altan, 2022) showed a new type of LSTM cell to handle chaotic time series prediction and reported that it outperforms traditional LSTM on several chaotic systems. Despite the domain knowledge studied in broad areas, developing new machine learning technologies for Chao’s purpose still needs more attention. In particular, the connection between chaotic prediction and conventional time-series forecasting is yet to be more closely integrated (Sahoo et al., 2019). There has been extensive literature on time-series forecasting as well  (PaperwithCodes, 2022), and sometimes these datasets hold chaotic properties as well, including using Transformer- and non-Transformer-based solutions for long-term time series forecasting (LTSF). For example,  (Mohammadi Farsani & Pazouki, 2020) showed the use of self-attention in the Transformer model for time-series prediction tasks that improved performance. Similarly,  (Huang et al., 2019) proposed the use of self-attention in a Recurrent Neural Network (RNN) architecture for time-series prediction, achieving improved performance on several datasets. However, these approaches still need more effective models for long-sequential input handling to consider the long history of the past. We have proposed a reservoir Transformer (RT) to reach this aim.
While reservoirs have found extensive application in time-series tasks, their performance often lags behind the current state-of-the-art outcomes (Gauthier, 2021). A concept of Reservoir Transformer  (Shen et al., 2020) was introduced by interspersing the reservoir layers into the regular Transformer layers and showing improvements in wall-clock compute time. However, such integration will reduce the feature extraction power of both reservoirs in the long-sequence and Transformer in short sequences. We do not combine the reservoir and Transformer linearly but in a non-linear way and receive significant improvements over all previous methods in time-series prediction, including chaotic prediction.
Additionally, there has been research on using ensembles of models for time-series prediction.(Feng et al., 2019) proposed an ensemble method that combines multiple hybrid models using wavelet transform for chaotic time-series prediction and achieved improved performance on several chaotic systems.(Tang et al., 2020) proposed an ensemble method that combines multiple deep-learning models for time-series prediction and showed improved performance on several datasets. In our work, we apply ensemble techniques to our Reservoir Transformer for the first time, contributing to further advancements in time-series prediction.

### v2

@@ Methods
TIME-LLM
GPT2TS
Methods
RT
TIME-LLM
GPT2TS
Metric MSE MAE MSE MAE MSE MAE Metric MSE MAE MSE MAE MSE MAE
ETTh1 96 0.361 0.386 0.362 0.392 0.376 0.397 Weather 96 0.146 0.196 0.147 0.201 0.162 0.212
192 0.396 0.412 0.398 0.418 0.416 0.418 192 0.193 0.236 0.189 0.234 0.204 0.248
336 0.427 0.422 0.4427 0.433 0.439 0.433 336 0.239 0.240 0.262 0.279 0.254 0.286
720 0.441 0.455 0.442 0.457 0.477 0.456 720 0.301 0.316 0.300 0.316 0.326 0.337
avg 0.408 0.423 0.408 0.423 0.465 0.455 avg 0.219 0.247 0.225 0.257 0.237 0.270
ETTh2 96 0.262 0.320 0.268 0.328 0.285 0.342 Electricity 96 0.130 0.222 0.131 0.224 0.139 0.238
192 0.331 0.373 0.329 0.375 0.354 0.389 192 0.154 0.243 0.152 0.241 0.153 0.251
336 0.358 0.403 0.368 0.409 0.373 0.407 336 0.158 0.245 0.160 0.248 0.169 0.266
720 0.369 0.411 0.372 0.420 0.406 0.441 720 0.190 0.294 0.192 0.298 0.206 0.297
avg 0.330 0.376 0.334 0.383 0.381 0.412 avg 0.158 0.251 0.158 0.252 0.167 0.263
ETTm1 96 0.272 0.336 0.272 0.334 0.292 0.346 Traffic 96 0.361 0.248 0.362 0.248 0.388 0.282
192 0.310 0.357 0.310 0.358 0.332 0.372 192 0.376 0.250 0.374 0.247 0.407 0.290
336 0.351 0.382 0.352 0.384 0.366 0.394 336 0.382 0.268 0.385 0.271 0.412 0.294
720 0.380 0.409 0.383 0.411 0.417 0.421 720 0.426 0.281 0.430 0.288 0.450 0.312
avg 0.328 0.371 0.329 0.372 0.388 0.403 avg 0.386 0.261 0.388 0.264 0.414 0.294
ETTm2 96 0.163 0.254 0.161 0.253 0.173 0.262 ILI 24 1.301 0.739 1.285 0.727 2.063 0.881
192 0.218 0.290 0.219 0.293 0.229 0.301 36 1.409 0.824 1.404 0.814 1.868 0.892
336 0.271 0.328 0.271 0.329 0.286 0.341 48 1.520 0.807 1.523 0.807 1.790 0.884
720 0.356 0.382 0.352 0.379 0.378 0.401 60 1.529 0.852 1.531 0.854 1.979 0.957
avg 0.252 0.316 0.251 0.313 0.284 0.339 avg 1.439 0.805 1.435 0.801 1.925 0.903
Avg. Training Time 3.150 Minutes 55.41 Minutes 14.76 Minutes Avg. Training Time 36.52 Minutes 510.2 Minutes 146.3 Minutes
Avg. Occupied GPU 2.033 GB 74.42 GB 16.31 GB Avg. Occupied GPU 2.110 GB 77.271 GB 17.680 GB
Table 3: Multivariate time-series forecasting (low chaos) comparison with pre-trained LLM model. ‘Avg. Training Time’ is the average training time for the model. ‘Avg. Occupied GPU’ is the average GPU memory usage for the model.
Method Air Quality DWV DGP DDF BTC AW TR
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE Acc. F-Score Acc. F-Score
Transformer 89.53 11.34 42.95 5.181 168.6K 353.3 1014 22.27 79.40K 232.1 0.928 0.940 0.649 0.934
RT (Ours) 70.58 9.040 6.181 1.893 28.79K 119.2 10.17 410.6 211.6 211.6 0.899 0.901 0.802 0.975
Table 4: Time series comparison (medium chaos) in MSE, MAE, Accuracy, and F-score with RT and baseline Transformer. ‘Acc.’ is the Accuracy.
Signature Feature Extraction: Table 6 shows the entropy of extracted features by baseline ($\mathbf{h}_{t}$), reservoir ($\mathbf{z}_{t}$), and RT ($\mathbf{f}_{t}$). Our RT model extracts more informative features than the baseline Transformer model. Using the Shannon entropy to measure the information amount, our reservoir output features $\mathbf{z}_{t}$ have higher entropy (more informative) than the embedding output features $\mathbf{h}_{t}$ in the baseline Transformer. Combining both $\mathbf{f}_{t}$ gives the highest entropy among them, demonstrating more information extraction with our RT. Details of the experiments are enclosed in the Appendix G.2.
Method RT (Ours) DLinear PatchTST
Horizons MGS ECG LA MGS ECG LA MGS ECG LA
50 0.173 0.301 0.313 0.204 0.321 0.356 0.179 0.297 0.310
100 0.185 0.314 0.331 0.225 0.343 0.370 0.186 0.315 0.313
500 0.206 0.341 0.359 0.240 0.371 0.389 0.209 0.347 0.360
1000 0.225 0.374 0.447 0.250 0.390 0.529 0.228 0.378 0.449
Table 5: MSE on three typical chaos datasets (high chaos).
Entropy
Datasets $\mathbf{h}_{t}$ $\mathbf{z}_{t}$ $\mathbf{f}_{t}$
ETTh1 1.011 1.966 2.331
ETTh2 0.993 1.892 2.015
ETTm1 0.816 2.056 2.831
ETTm2 0.599 1.248 2.131
Table 6: Feature entropy.
4.1 Ablation Analysis
Mackey Glass Series.
ECG Signal.
Air Quality.
Figure 2: The figures represent chaotic dataset examples. The cyan line is ground truth; the red line is the prediction using RT; and the blue line is the prediction using Transformer.
Output Examples: Figure 1 (Right) and Figure 2 plots the examples from three datasets. We show that our RT method output is more close to the ground truth than the Transformer baseline.
Figure 3: MSE vs. reservoir numbers.
Figure 4: RT outperforms baselines with small parameter size and loss in training.
Impact of Reservoir Count on Performance: Figure 3 illustrates the relationship between the number of ensembling reservoirs and system performance. Increasing the number of reservoirs improves the performance, leading to a reduction in loss until the loss converges for $10$ reservoirs.
@@ 4.1 Ablation Analysis
Model Parameter Size versus Loss: Conventional models require a large number of training parameters for long historical sequences. In contrast, RT only train a vector representation (i.e. readout) for the next time step without explicit conditioning on the entire history. Figure 4 illustrates that RT provides a better trade-off between the number of training parameters and loss.
Initialization ETTh1 ETTh2 ETTm1 ETTm2 Weather Traffic
NG G NG G NG G NG G NG G NG G
Random 0.543 0.402 0.435 0.342 0.447 0.328 0.301 0.228 0.334 0.198 0.424 0.380
Zero 0.429 0.396 0.368 0.331 0.348 0.310 0.236 0.218 0.215 0.193 0.380 0.376
Constant 0.673 0.405 0.420 0.333 0.516 0.319 0.265 0.225 0.204 0.196 0.424 0.378
Normal 0.529 0.401 0.445 0.335 0.435 0.332 0.274 0.236 0.242 0.198 0.452 0.379
Uniform 0.432 0.399 0.395 0.331 0.401 0.314 0.232 0.219 0.200 0.195 0.401 0.375
[…]
Table 7: MSE on different initialization using single reservoir (NG) and group reservoirs (G). Group reservoir eliminates the prediction error sensitivity on the initialization.
Model Memory (Cache) Footprint Context Length Model Time Complexity Seq. Ops.
Transformer-XL $(d_{key}+d_{value})\cdot H\cdot k\cdot ly$ $k\cdot ly$ Transformer $O(T^{2}\cdot d_{\epsilon})$ $O(1)$
Compressive Transformer $d_{\epsilon}\cdot(c+T)\cdot ly$ $(c\cdot r+k)\cdot ly$ PatchTST $O((T/P)^{2}\cdot d_{\epsilon})$ $O(1)$
Memorizing Transformers $(d_{key}+d_{value})\cdot H\cdot k\cdot T$ T Informer $O(T\cdot log(T)\cdot d_{\epsilon})$ $O(1)$
RMT $d_{\epsilon}\cdot p\cdot ly\cdot 2$ T Autoformer $O(T\cdot log(T)\cdot d_{\epsilon})$ $O(1)$
AutoCompressors $d_{\epsilon}\cdot p\cdot(v+1)\cdot ly$ T Reformer $O(T\cdot log(T)\cdot d_{\epsilon})$ $O(1)$
Infini-Transformers $d_{key}\cdot(d_{value}+1)\cdot H\cdot ly$ T RNN $O(T\cdot d_{\epsilon}^{2})$ $O(T)$
Reservoir $N_{r}^{2}$ T Reservoir $O(T\cdot N_{r}^{2})$ $O(T)$
RT (Ours) $N_{r}^{2}\cdot L$ T RT (Ours) $O(T\cdot k^{2}\cdot d_{\epsilon})$ $O(T)$
Table 8: Comparison of memory (cache) footprint and time complexity.
Initialization Comparison: Table 7 shows how different initialization affects the performance for each of the datasets. We compute the the $p$-value to see whether the variation in performance due to different initialization is significant when comparing a single reservoir (Non-Group denoted with ‘NG’) with an group reservoir (Group denoted with ‘G’). We use the combined mean for each dataset as the population mean to compute the $p$-values. We observe significant differences for all the datasets for only the Non-Group experiments.
Complexity Comparison: Table 8 shows the memory (cache) footprint and time complexity of different models. ‘Seq. Ops.’ is the ‘Sequential Operations’, and the notations are as follows. $N_{r}$ is the reservoir state; $L$ is the number of reservoirs in the group reservoir; $k$ is the short-term context window length; $T$ is the long-term context total input length; $ly$ is the number of layers; $H$ is the number of attention heads; $c$ is the Compressive Transformer memory size; $r$ is the compression ratio; $p$ is the number of soft-prompt summary vectors; $v$ is the summary vector accumulation steps; $s$ is the kernel size; $P$ is the patch size; $d_{\epsilon}$ is the embedding model dimension; $d_{key}$ is the dimension of keys in attention; and $d_{value}$ is the dimension of values in attention. RT maintains a constant memory complexity of $N_{r}^{2}\cdot L$, and the time complexity is linear regarding total input length. Additionally, RT updates its memory with each new input, resulting in retrieval and update occurring once per input. Other methods like Transformer-XL (Dai et al., 2019), Compressive Transformer (Rae et al., 2019), Memorizing Transformers (Wu et al., 2022b), RMT (Bulatov et al., 2022), and AutoCompressors (Chevalier et al., 2023) do not have constant memory usage; it depends on the total input segment length $N$ and still exhibits a large context-memory complexity. Furthermore, RMT (Bulatov et al., 2022) and AutoCompressors (Chevalier et al., 2023) allow for a potentially infinite context length since they compress the input into summary vectors and then pass them as extra soft-prompt inputs for the subsequent segments. However, in practice, the success of those techniques highly depends on the size of soft-prompt vectors (Munkhdalai et al., 2024).
@@ 5 Related Work
Time-series prediction using machine learning approaches has been investigated in the last decade using CNN, WNN, FNN, LSTM, etc. (Ramadevi and Bingi, 2022; Sangiorgio and Dercole, 2020; Karasu and Altan, 2022; Sahoo et al., 2019; PaperwithCodes, 2022) on datasets with chaotic properties. Recently, Transformer-based solutions have shown success for long-term time series forecasting (LTSF) (Mohammadi Farsani and Pazouki, 2020; Huang et al., 2019). Recently, researchers have also used various methods to improve long-term time series forecasting  (Zhou et al., 2022, 2021; Wu et al., 2021). Nonetheless, the Transformer is expensive in terms of quadratic time and space complexity of the input length. Important works have been proposed in improving Transformer architectures and their complexity in long context in various ways, however, compressing the long input sequence into fixed length often misses the dependency of the temporal events.
Researchers have also shown how using pre-trained LLMs for time-series data improves performance compared to training from scratch (Jin et al., 2023; Chang et al., 2024; Huang, 2024; Munkhdalai et al., 2024). de Zarzà et al. (2023) used a combination of LLMs (LLaMA-2 and Zephyr-7b-$\alpha$) along with Multimodal LLM (LLaVA) for time-series forecasting. However, there is still a restriction on the input length due to the complexity of the neural network models. Note that we will experimentally compare with (Munkhdalai et al., 2024) once their code releases.
Reservoir networks with their simplified training and superior performance for time-series forecasting compared to the traditional models, including time-series kernels, feed-forward networks, and RNN models, has been well documented. Researchers have also compared different types of reservoir networks for chaotic time-series data (Bo et al., 2020). However, the reservoir network fails to effectively handle long context time series data within fixed-size reservoir (Lukoševičius and Jaeger, 2009; Kim and King, 2020; Mansoor et al., 2021; Bianchi et al., 2020). We have introduced reservoir computing to solve the input length problem. A concept of reservoir Transformer  (Shen et al., 2020; Gauthier et al., 2021) was introduced by interspersing the reservoir layers into the regular Transformer layers and showing improvements in wall-clock compute time. These models improve the training efficiency by using parallelization of time-series methods  (Bollt, 2021). However, such integration will reduce the feature extraction power of both reservoirs in the long sequences and Transformer in short sequences. We do not combine the reservoir and Transformer linearly but in a non-linear way and receive significant improvements over all previous methods in time-series prediction, including chaotic prediction. Additionally, there has been research on using ensemble of models for time-series prediction. Researchers have proposed ensemble methods (Feng et al., 2019) that combine multiple hybrid models using wavelet transform for chaotic time-series prediction and achieved improved performance on several chaotic systems.(Tang et al., 2020; Oliveira and Torgo, 2015; Bertsimas and Boussioux, 2023) proposed an ensemble method that combines multiple deep-learning models for time-series prediction and showed improved performance on several datasets. In our work, we propose a novel method to ensemble multiple reservoir networks and combine them with Transformer architecture, contributing to improved performance in time-series prediction.